OpenStudy (anonymous):

30% of what number is 42?

OpenStudy (mathlegend):

Try... %/100 = is/of and cross multiply @lexusleamon

OpenStudy (anonymous):

ok t5hanxs

OpenStudy (anonymous):

30x=42 x= 42/30 x = 1.4

OpenStudy (anonymous):

tk

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

0.30n=42 42/0.30=n n=140